OpenStudy (anonymous):

The two shorter sides of an obtuse triangle are 20 and 30. What is the least whole number length possible for the third side? a. 35 b. 45 c. 37 d. 47 Thanks! x

OpenStudy (jojo4eva):

b.45

OpenStudy (anonymous):

Thanks so much! Would you mind explaining? :) Just so I know how to solve in the future

OpenStudy (jojo4eva):

ok so basically i used the a^2+b^2=c^2. 20^2=400 and 30^2=900 900+400=1300. OMg i messed up im sorry the answer is c.37 because c=37 so 37*37=1369 which is closest to 1300

OpenStudy (anonymous):

Ok, so I was going to use the Pythagorean theorem to answer this but it doesn't say that it's a right triangle. It says it's an obtuse triangle. So how can you use a²+b²=c²?

OpenStudy (anonymous):

Or, you can simply say all triangles' angles have to equal up to 180 degrees right? And the definition of Obtuse means that the angle is larger than 90 degrees. So any number you get that's greater than 90 is considered an Obtuse angle

OpenStudy (jojo4eva):

u kno what you are exactly right wow im sorry i messed you up. @phi can u plz help this person out i dont want to spam either but they need help. Im sorry @ajm i wasnt paying attention

OpenStudy (anonymous):

Haha that's ok! Thanks though:)

OpenStudy (anonymous):

@K-K-K-Kiaa yeah, but it's dealing with side lengths, not angles...

OpenStudy (anonymous):

Ahhh okay sorry, It can't have a leg longer than the sum of its other two sides or shorter than the difference.

OpenStudy (anonymous):

Yeah… I tried that too. As far as I can tell, all of them work like that

OpenStudy (anonymous):

No, look again. It says "the least whole number length"

OpenStudy (anonymous):

So it would be a. 35 just bc it's the smallest number?

OpenStudy (anonymous):

Correct, you are!

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

No problem!